OpenStudy (anonymous):

y = –x^2 – 6x – 5 , find the zeros of the functions w/steps please.

hero (hero):

To find the zeroes of the function, 1. Set y = 0: y = -x^2-6x-5 0 = -x^2-6x-5 2. factor a value common to each term. In this case -1 is the common factor: 0 = -1(x^2 + 6x + 5) 3. Divide both sides by -1 0 = x^2 + 6x + 5 4. Factor as usual by finding two numbers that multply to get 5, yet add to get 6: The two numbers are 5 and 1. Use those two numbers to split the middle term then continue factoring: 0 = x^2 + 5x + 1x + 5 0 = x(x+5)+1(x+5) 0 = (x+5)(x+1) 5. Use the zero product property: x + 5 = 0 x + 1 = 0 Thus the two zeroes are x = -1,-5

OpenStudy (anonymous):

Got it. Thanks Hero!

hero (hero):

yw =]